# quality of service 5G fifth generation

### v2

@@ Abstract
Resource allocation, Network slicing, O-RAN, Variational Autoencoder, Contrastive Learning
I INTRODUCTION
Wireless communication has evolved from a mere convenience to an absolute necessity, driven by an insatiable demand for faster and more dependable connectivity between various quality of service (QoS) requirements [1]. The emergence of fifth generation (5G) technology represents a significant leap, offering the potential to support a wide range of applications, including augmented reality and autonomous vehicles [2]. However, the pursuit of wireless excellence does not conclude with 5G; it marks a stepping stone towards the next frontier— sixth generation (6G), the sixth generation of wireless technology. Poised on the horizon, 6G promises to redefine the boundaries of connectivity again.
Emerging concepts within 6G, such as extreme Multiple-Input Multiple-Output and deep Receiver, promise to propel us further into uncharted technological realms, offering unprecedented data rates and levels of connectivity [3]. These advancements can potentially revolutionize our modes of communication and reshape the fabric of our interconnected world. Furthermore, the application of various machine learning (ML) techniques to automate resource allocation within the medium access control (MAC) layer and scheduler represents a burgeoning area of research within the 6G landscape [4]. These techniques aim to inject unparalleled intelligence into network management and optimization, laying the foundation for networks capable of dynamically adapting to user demands and environmental conditions. In this constantly evolving landscape, the roles of Open Radio Access Network (O-RAN) [5] and network slicing [6], which have reached remarkable milestones in 5G, continue to be pivotal. As we explore the possibilities offered by 6G and its transformative capabilities, it becomes increasingly clear that the evolution of wireless communication is an ongoing journey, establishing a robust foundation for the next era of wireless excellence.
Resource allocation involves carefully distributing limited network resources among numerous users, applications, or services [7]. In 5G, the main goal is to support a wide range of services, including Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and massive Machine Type Communications (mMTC) [8]. Each service requires specific resource sets, including bandwidth, data rate, latency, power consumption, complexity, and reliability. Resource allocation within the 5G landscape needs to be flexible to meet these diverse demands while efficiently using network resources.
Network slicing is an innovative concept that customizes virtual networks (slices) within a unified physical infrastructure to address specific service types and use cases [9]. In 5G, it efficiently allocates resources for eMBB and URLLC [10] services. This significance extends into the anticipated 6G era, where network slicing remains instrumental, facilitating resource allocation for diverse applications, including holographic communication and tactile internet [11, 12]. Its inherent scalability ensures adaptability to evolving workloads and user expectations, solidifying its role as a critical technology in the transition from 5G to 6G.
This concept aligns with the architecture of O-RAN. O-RAN is based on open interfaces, open hardware, and open software and aims to provide a more flexible and cost-effective solution for wireless communication [7]. Implementing network slicing in O-RAN optimizes resource allocation and promotes multi-vendor interoperability and innovation.
The core of O-RAN’s architecture consists of two fundamental components: Radio Units, responsible for physical layer functions, and Distributed Units, which manage baseband processing and control. These components work together, with the Central Unit (CU) serving as a centralized coordinator for network resources, facilitating dynamic resource allocation [13, 5]. The CU further comprises two distinct elements: the control plane (CP) and the user plane (UP). Adding to O-RAN’s adaptability and innovation are xApps (cross-applications), capable of spanning multiple network domains to optimize resource allocation for various services. These network segments include Fronthaul, Backhaul, and Midhaul, facilitating data transmission between components. Overseeing resource management and orchestration tasks is the Management and Orchestration (MANO) layer, ensuring the network’s efficiency. Network Functions Virtualization (NFV) decouples software from hardware, enhancing scalability. Furthermore, the Radio Access Network Intelligent Controller (RIC) plays a pivotal role in intelligent resource control, enabling dynamic allocation to meet diverse service requirements [5]. These components collectively define O-RAN’s transformative architecture, promoting openness, adaptability, and efficient resource utilization in modern wireless networks.
It is worth noting that O-RAN comprises two types of RIC: near-real-time (near-RT) RIC, which manages Radio Access Network (RAN) control and optimization, and non-real-time (non-RT) RIC. Within the near-RT RIC, xAPPs facilitate RAN control and optimization, while inside the non-RT RIC, rAPPs support it by providing guidance and employing ML models. Furthermore, both xAPPs and rAPPs can be implemented by third-party entities, allowing for a diverse ecosystem of applications within the O-RAN architecture [14].
Using xAPPs leverages O-RAN network slicing to use resources efficiently and facilitate collaboration across services. However, this also introduces complexities in network management, as each xAPP needs to adjust the different QoS requirements of multiple services, posing a coordination challenge. Furthermore, cellular operators face the difficulty of establishing consistent connectivity for devices with varying QoS demands [15, 16]. Therefore, this study addresses the resource allocation problem within O-RAN using network slicing by considering the xAPPs. We propose a novel resource allocation method to effectively coordinate the numerous independent xAPPs to address network slicing and resource allocation challenges.
The use of ML algorithms in wireless networks has increased due to the growing complexity of these networks and the challenges of creating custom optimization models. Traditional approaches, such as convex/non-convex optimization, are struggling with the increasing complexity stemming from changes in network architecture, evolving performance requirements, and the growing number of devices. Current methods for network resource allocation that rely on learning-based techniques, like deep learning (DL) and deep reinforcement learning (DRL), often overlook the ability to learn resource distributions. Typically, these methods involve designing an agent for a specific task and training it from scratch, which reduces exploration efficiency and extends convergence times. Moreover, they require many training samples, leading to prolonged exploration phases.
The limited adaptability of traditional ML methods to new scenarios motivates us to explore an alternative approach capable of better generalization, more efficient data utilization, and resource distribution learning. To address these objectives, we propose a novel methodology that uses a self-supervised learning technique, specifically contrastive learning [17], in combination with a variational autoencoder (VAE) model, which is a type of generative model [18]. These models excel in handling uncertainty by learning distributions, offering a distribution for predictions instead of a singular deterministic output.
Moreover, their skill in handling incomplete or missing data allows them to fill in missing values based on learned patterns. VAEs also excel in semi-supervised learning setups, effectively using labeled and unlabeled data to improve predictive accuracy. Their generative abilities enable the creation of new data instances that resemble the training set, which is helpful in scenarios requiring synthetic data generation or increased resilience against noise and outliers.
Furthermore, contrastive learning involves comparing different examples to acquire informative data representations. On the other hand, a VAE consists of an encoder and decoder that work together to learn the underlying probability distribution of data. By integrating contrastive learning with VAE, we can address the challenge of needing many training samples in ML. This approach can improve performance metrics, leading to faster convergence and better results while reducing the time and effort required for algorithm design and training.
In our proposed solution, x-Apps help to implement advanced resource allocation algorithms within the O-RAN framework. The algorithm is deployed as an x-App on the near-real-time RIC, enabling efficient and dynamic resource allocation adjustment based on real-time network conditions and user demands.
I-A Contribution
This paper introduces a system model in the O-RAN architecture that provides support for two primary services, namely eMBB and URLLC. It focuses on resource allocation, with a particular emphasis on network slicing. The primary contributions of our paper can be summarized as follows:
  • Our research aims to address the resource allocation problem in the O-RAN architecture, explicitly focusing on optimizing weighted throughput by allocating resources such as transmission power and Physical Resource Blocks. We consider the unique QoS requirements and service priorities of two key 5G services: eMBB and URLLC. Compared to existing joint resource allocation schemes, our xAPP-based architecture enhances network management flexibility.
  • Our approach to optimizing resource allocation involves two main steps. Firstly, we simplify the problem by framing it as an optimization problem. Then, we introduce a novel machine learning algorithm that combines supervised learning with a VAE and unsupervised learning with contrastive loss to optimize resource allocation in O-RAN. It significantly improves the model’s generalization ability, making it a more practical solution for real-world applications.
  • While explicitly guaranteeing a global optimal solution is difficult, our method aims to achieve near-optimal solutions through evaluation with multiple benchmarks and significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems. The method balances computational efficiency and scalability, making it suitable for real-time resource allocation in dynamic O-RAN environments. It is not limited to O-RAN and can be applied to future multi-vendor O-RAN architectures, leveraging semi-supervised generative model principles developed by our team.
I-B Structure of Paper
The paper is structured as follows: Section II reviews related works. In Section III, we introduce the system model for the resource allocation problem in the O-RAN architecture and formally define the resource allocation problem as an optimization problem. Section IV provides a detailed description of the proposed resource allocation algorithm. Section V presents numerical results used to evaluate the algorithm’s performance. Section VI compares the computational cost of employed algorithms. Finally, Section VII offers our conclusion.
II Related Works
As mentioned in Section I, cellular operators face significant challenges in ensuring reliable connections for devices with diverse QoS requirements. This challenge becomes more prominent in 5G and 6G networks due to their increased complexity and stringent QoS demands [16]. Recently, a growing focus has been on addressing the network slicing problem in cellular networks [16, 19, 20, 21, 22].
Feng et al. [16] proposed a method that employs network slicing and scheduling to allocate resources based on QoS requirements. Another study presented a resource allocation approach based on the M/M/1 queuing system for RAN slicing [19]. Nevertheless, practical scenarios often involve bursty traffic, overly optimistic, rendering the M/M/1 assumption of a Poisson distribution of packet arrivals at each gNodeB, which serves the eMBB and URLLC services. Consequently, the gNodeB cannot be considered an M/M/1 system, and its queue cannot be accurately modeled as a continuous-time Markov process. Additionally, in a study by Kazemifard et al., [20], an O-RAN system was investigated, assuming that the core network and RAN are co-located in the same data center. The authors conducted a theoretical analysis and employed a mathematical model to address minimum delay problems efficiently.
Various approaches have been proposed in the literature to address resource allocation challenges in wireless networks, including optimization techniques, game theory, and ML algorithms [23]. ML algorithms, in particular, have found application in power control and resource allocation within wireless networks.
For instance, the SAMUS framework was employed by Bektas et al. [24] to forecast incoming traffic, enabling an exploration of various network slicing operations and a discussion of the trade-off between the data rate of eMBB and the latency of URLLC. Li et al. [25] introduced a method that utilizes DRL for proactive resource allocation to mobile UE, considering both time and frequency domains while accounting for mobility. Wu et al. [26] proposed a dynamic resource allocation method for radio and computation resources through RAN slicing, including a defined learning scheme with constraints.
Further, Ayala et al. [27] presented a method for the joint allocation of virtualized RAN’s radio and computation resources using a DRL approach based on an actor-critic network. Wang et al. [28] introduced a DRL-based self-play approach that can be integrated into RAN architecture to efficiently manage RU and DU resources, leading to cost-effective RAN operations. In another approach, Ndikumana et al. [29] combined federated learning and Deep Q-Learning with a reward function to reduce offloading, fronthaul routing, and computation delays in O-RAN.
Similarly, Samarakoon et al. [30] introduced a federated learning algorithm to optimize power transmission and radio resource allocation in vehicular networks. Additionally, Elsayed et al. [31] proposed a reinforcement learning-based method for joint power and radio resource allocation in 5G networks.
III System Model and Problem Formulation
In our study, we address the resource allocation problem using network slicing within O-RAN architecture, focusing on two primary slices: the URLLC slice and the eMBB slice. The URLLC slices ensure ultra-low latency and high reliability for critical applications such as autonomous driving and remote surgery. In contrast, eMBB slices focus on high data rates for mobile broadband. Both slices have different QoS requirements for network resource allocation [10].
[…]

### v3

@@ Abstract
Generative AI, Variational Autoencoder, Contrastive Learning, Open Radio Access Network (O-RAN) Slicing, Resource Allocation
1 Introduction
Wireless communication has evolved from a convenience to a necessity, driven by the increasing demand for faster and more reliable connectivity across diverse Quality of Service (QoS) requirements [1]. The advent of Fifth Generation (5G) marks a significant leap, enabling applications such as augmented reality and autonomous vehicles [2], while laying the foundation for Sixth Generation (6G), which aims to further redefine global connectivity.
To meet the growing complexity of future networks, 3GPP begins integrating Artificial Intelligence (AI) into core Radio Access Network (RAN) functions such as beam management, positioning, and Channel State Information (CSI) feedback, initiating a shift towards intelligent and adaptive systems. As technologies like extreme Multiple-Input Multiple-Output (MIMO), deep receivers, and deep schedulers emerge, AI becomes essential to handle real-time data-driven tasks such as Physical Resource Blocks allocation and power control in both uplink and downlink [3, 4, 5].
Network slicing and O-RAN are critical enablers in both 5G and 6G, offering the flexibility to support services like Enhanced Mobile Broadband (eMBB), Ultra-Reliable Low Latency Communications (URLLC), and Massive Machine-Type Communications (mMTC), each with distinct performance demands [6, 4]. In 6G, network slicing will extend to advanced use cases such as holographic communication and tactile Internet, ensuring scalable and adaptable resource management [7, 8]. At the same time, O-RAN—with its open interfaces, disaggregation of hardware and software—provides a modular, cost-effective platform for embedding AI-based control and enabling multi-vendor interoperability [9].
These architectures are foundational for realizing sophisticated AI-driven resource allocation strategies. As network demands grow and QoS requirements diversify, highly adaptive intelligent solutions for dynamic scheduling, power control, and PRB management become increasingly paramount. Effectively integrating such advanced AI within flexible frameworks such as network slicing and O-RAN is crucial to ensure future wireless systems are not only efficient and responsive but also proactively service-aware, making sophisticated AI capabilities indispensable for unlocking the full potential of 6G [10, 11].
The inherent complexity and dynamism of these emerging wireless environments have spurred the adoption of AI algorithms, as conventional optimization methods often lack the agility to cope with rapidly evolving network states and diverse performance criteria. Although learning-based paradigms, such as Deep Learning (DL), have demonstrated potential for specific RAN optimization tasks, their application to multifaceted resource allocation in O-RAN slicing often encounters significant hurdles. These typically include the need for extensive, task-specific agent engineering, computationally intensive training regimens that usually start from scratch, and practical limitations related to sample inefficiency and suboptimal generalization to novel network conditions or previously unseen service demands. Such drawbacks critically hinder the development of truly autonomous systems capable of generating robust, adaptive, and near-optimal resource management strategies in real time.
The previously discussed limitations of conventional AI methods in achieving adaptable and data-efficient resource allocation for dynamic networks underscore the need for novel Generative Artificial Intelligence (GAI) paradigms. To this end, we propose an innovative GAI framework, termed Generative Semi-Supervised Variational Autoencoder (VAE)-Contrastive Learning (SS-VAE), designed explicitly for intelligent O-RAN network slicing. This framework is realized through a unified architecture that intrinsically integrates the distributional learning and generative capabilities of a VAE [12] with the potent representation learning power of contrastive learning principles [13]. Trained end-to-end using a semi-supervised strategy, our SS-VAE model is designed to learn complex resource distributions and generate optimized allocation decisions with superior generalization and data efficiency.
The synergistic design of our Generative SS-VAE framework offers distinct advantages for GAI-driven resource optimization in O-RAN. The VAE core excels at modeling the underlying probability distributions of network states and resource utilization, which is crucial for managing the inherent uncertainty and can be leveraged to handle incomplete data by inferring missing values from learned patterns. Its generative capability, central to our GAI approach, allows the model to proactively construct or refine resource allocation strategies, finely tuned to diverse QoS demands within O-RAN slices. This is powerfully augmented by the integrated contrastive learning mechanism, which compels the model to learn highly informative and robust latent representations from both limited labeled data and abundant unlabeled data. Such a focus on representation quality, integral to our semi-supervised methodology, significantly enhances sample efficiency and fortifies the model’s ability to generalize to unseen network conditions and resist noise or outliers, making it exceptionally suited for the dynamic O-RAN environment.
Our proposed SS-VAE algorithm is designed for practical implementation within the O-RAN architecture. It is envisioned for deployment as an eXtended Application (xApp) operating on the near-real-time Radio Access Network Intelligent Controller (RIC). This strategic placement enables the SS-VAE xApp to dynamically adjust resource allocation by leveraging real-time network telemetry and user demand information, thereby facilitating efficient, adaptive, and QoS-aware management of network slices.
1.1 Contribution
This paper makes significant contributions to the field of intelligent resource allocation and network slicing within the O-RAN architecture, explicitly targeting the distinct demands of eMBB and URLLC services. We introduce and rigorously evaluate a novel GAI driven solution. The primary contributions of this work are as follows.
  • We present a detailed system model and formulate the resource allocation problem in O-RAN environments, aiming to optimize weighted throughput via efficient assignment of transmission power and PRBs in the downlink. This formulation explicitly considers the unique QoS requirements and service priorities of eMBB and URLLC services. It is designed for implementation within an xApp-based architecture to enhance the flexibility of AI-driven network management.
  • We propose, design and implement a novel GAI framework to optimize resource allocation in O-RAN. This framework employs a unified deep learning architecture that combines a VAE—which captures complex data patterns and can generate resource allocation solutions—with contrastive learning to learn robust and meaningful feature representations. The entire model is trained end-to-end using a semi-supervised approach, which improves both data efficiency and overall performance. This enables effective generalization across various O-RAN environment scenarios. This makes the framework highly adaptable and practical for real-world, dynamic O-RAN systems.
  • We demonstrate through extensive simulations and comparative analysis against multiple benchmarks—including the exhaustive search algorithm (ESA) and deep reinforcement learning—that our proposed AI-driven approach achieves near-optimal resource allocation performance with substantially lower computational complexity than state-of-the-art methods. The results validate the robustness, computational efficiency, and scalability of our AI solution, affirming its suitability for real-time resource management in dynamic O-RAN systems. Furthermore, the principles of our semi-supervised generative model pave the way for its application in future multi-vendor O-RAN architectures and evolving service paradigms.
1.2 Structure of the Paper
The paper is structured as follows: Section 2 provides a brief overview of the O-RAN architecture and its components relevant to intelligent control. Section 3 reviews related works in resource allocation, particularly focusing on AI-based approaches in O-RAN. In Section 4, we introduce the system model for the resource allocation challenge in the O-RAN architecture and formally define the problem as an optimization task. Section 5 elaborates on our novel Generative AI-driven resource allocation framework, detailing its underlying model architecture and training methodology. Section 6 presents the numerical results from our performance evaluation, comparing our approach against benchmark algorithms. Section 7 provides an analysis of the computational complexity of the employed algorithms. Finally, Section 8 offers concluding remarks and outlines potential future research directions. For clarity, the following section lists the main acronyms used throughout the paper.
List of Acronyms and Abbreviations
5G
Fifth Generation
6G
Sixth Generation
AI
Artificial Intelligence
AMF
Access and Mobility Management Function
BS
Base Station
C-RAN
Cloud Radio Access Network
CP
Control Plane
CSI
Channel State Information
CU
Central Unit
DL
Deep Learning
DM
Diffusion Model
DNN
Deep Neural Network
DQN
Deep Q-Network
DRL
Deep Reinforcement Learning
DU
Distributed Unit
eMBB
Enhanced Mobile Broadband
ESA
Exhaustive Search Algorithm
FDRL
Federated Deep Reinforcement Learning
FL
Federated Learning
GAI
Generative Artificial Intelligence
GAN
Generative Adversarial Network
KL
Kullback-Leibler
LLM
Large Language Model
MAC
Medium Access Control
MAE
Mean Absolute Error
MANO
Management and Orchestration
MARL
Multi-Agent Reinforcement Learning
MDP
Markov Decision Process
MIMO
Multiple-Input Multiple-Output
ML
Machine Learning
mMTC
Massive Machine-Type Communications
MSE
Mean Squared Error
near-RT
Near-Real-Time
NFV
Network Functions Virtualization
NLOS
Non-Line-of-Sight
non-RT
Non-Real-Time
O-RAN
Open Radio Access Network
OPEX
Operational Expenditures
PDCP
Packet Data Convergence Protocol
PRB
Physical Resource Block
QoS
Quality of Service
RAN
Radio Access Network
RB
Resource Block
RIC
Radio Access Network Intelligent Controller
RL
Reinforcement Learning
RU
Radio Unit
SDAP
Service Data Adaptation Protocol
SINR
Signal to Interference & Noise Ratio
TL
Transfer Learning
UE
User Equipment
UP
User Plane
UPF
User Plane Function
URLLC
Ultra-Reliable Low Latency Communications
VAE
Variational Autoencoder
vRAN
Virtualized Radio Access Network
xApp
eXtended Application
2 Background
Figure 1: O-RAN network slicing
The O-RAN Alliance has created an innovative RAN framework designed to enable an open, intelligent, virtualized, and interoperable RAN, which is crucial for affordable, next-generation wireless networks [14]. The O-RAN architecture combines Cloud Radio Access Network (C-RAN) and Virtualized Radio Access Network (vRAN) to create open, interoperable, and flexible mobile networks that support both 5G and future 6G requirements. The core components of O-RAN include the Open Radio Unit (O-RU), the Open Distributed Unit (O-DU), and the Open Centralized Unit (O-CU), each responsible for specific layers of the protocol stack. The O-RU handles radio frequency (RF) and lower physical (PHY) functions. The O-DU processes the medium access control (MAC), radio link control (RLC), and higher PHY layers. Meanwhile, the O-CU manages the Packet Data Convergence Protocol (PDCP) and Service Data Adaptation Protocol (SDAP) and is further divided into the Control Plane (CP) and User Plane (UP) for efficient signaling and traffic routing [15, 14]. Beyond these disaggregated data plane components, key O-RAN architectural pillars include Orchestration and Automation, the O-Cloud, and notably, the RAN RIC. The RIC is pivotal in enabling intelligent network operations, facilitating near-real-time and non-real-time optimization and learning-based control over the RAN.
In the subsequent section, we provide a brief overview of the key methods and characteristics employed in the O-RAN system, which improve its adaptability and efficiency.
2.1 Network Slicing in O-RAN
A cornerstone of 5G/6G models, network slicing facilitates the creation of customized virtual networks on demand that operate on shared infrastructure. This end-to-end capability is achieved by logically separating and managing resources across three key domains: the RAN, the core, and the transport network. In the RAN domain, slicing involves partitioning radio resources, such as PRBs, and instantiating dedicated virtual network functions for different services, including MAC/RLC in the O-DU and PDCP/SDAP in the O-CU, as illustrated in Figure 1. In the core network, key nodes, such as the User Plane Function (UPF) and Access and Mobility Management Function (AMF), are virtualized and isolated to create slices that accommodate specific service-level agreements. The transport network is sliced to create dedicated data pathways, ensuring that the performance and QoS for these diverse connections are guaranteed. The O-RAN framework, with its inherent intelligence and virtualization, serves as a crucial enabler for sophisticated RAN slicing, which is indispensable to perform comprehensive end-to-end network slicing [16, 17].
2.2 Radio Intelligent Controller (RIC)
The RIC is functionally divided into two layers: the near-real-time (near-RT) and the non-real-time (non-RT) RIC. The near-RT RIC, designed for control loops operating within 10 milliseconds to 1 second latency, hosts modular xApps that enable fast and responsive control over RAN functions. These xApps handle tasks such as interference mitigation, handover control, and dynamic resource allocation, interacting with the RAN in real time. They are fundamental to enabling service-level customization, particularly through the use of network slicing. In contrast, the non-RT RIC, typically situated within the broader Orchestration and Automation framework, manages operations with latency requirements exceeding 1 second. It supports functions like policy management, network analytics, and AI model training by hosting rApps. These applications analyze historical data, orchestrate the training of AI models, and generate strategic policies or guidance for the near-RT RIC and its xApps [18, 19].
Together, the non-RT RIC and near-RT RIC establish a hierarchical intelligent control loop. Typically, AI models are trained or designed in the non-RT RIC using historical data and network policies and then deployed or used by xApps in the near-RT RIC for inference and real-time action. This synergy enables predictive and adaptive RAN behavior, significantly enhancing automation in dynamic network environments. Both xApps and rApps are designed to be modular and leverage open interfaces, which encourages third-party innovation and accelerates the deployment of AI-driven functionalities, thereby promoting multi-vendor interoperability within the O-RAN ecosystem.
2.3 Artificial Intelligence (AI) in O-RAN
The integration of AI is a foundational tenet of the O-RAN architecture, designed to introduce intelligence and autonomy into network operations. This approach is crucial for enhancing QoS through dynamic optimization and reducing Operational Expenditures (OPEX) through automated management. The application of AI is envisioned across a spectrum of RAN use cases, from real-time resource allocation and traffic steering to anomaly detection and cybersecurity.
To address these challenges, the O-RAN framework is equipped to leverage a diverse range of AI methodologies. Supervised learning enables predictive capabilities, such as traffic forecasting, by training on labeled historical data. In contrast, unsupervised learning identifies latent patterns in unlabeled data for tasks such as user clustering or threat detection. For autonomous control, Reinforcement Learning (RL) and its deep-learning variant ( Deep Reinforcement Learning (DRL)) allow agents to derive optimal policies through direct interaction with the network environment. To address privacy concerns in multi-agent environments, Federated Learning (FL) facilitates collaborative model training on distributed data without centralizing sensitive information. Further advancing these capabilities, GAI models like VAEs can learn data distributions to generate synthetic training data or even propose new resource management solutions directly. These core techniques can be augmented by Large Language Models, which provides a semantic understanding of unstructured data to enhance the overall intelligence of the system’s decision-making.
These AI functionalities are operationally hosted within the O-RAN Intelligent Controllers (RIC). The non-RT RIC typically manages functions with relaxed latency constraints, such as offline model training or acting as the central server in a Federated Learning setup. Conversely, the near-RT RIC executes tasks requiring rapid response, including model inference for sub-second control loops and the distributed training aspects of RL and FL. This architectural division allows O-RAN to effectively support the entire AI lifecycle, from model creation in the non-RT RIC to real-time deployment and execution in the near-RT RIC.
3 Related Works
As highlighted in Section 1, ensuring reliable connectivity for devices with diverse QoS requirements presents significant challenges for cellular operators, especially in the complex and demanding landscape of 5G and emerging 6G networks [20]. As a result, network slicing has become a key focus of research in cellular networks. With the increasing complexity and diversity of services, methodologies for resource allocation have undergone significant evolution. They have shifted from traditional mathematical optimization and heuristics to more adaptive and intelligent AI-driven approaches [20, 21, 22, 23, 24].
3.1 Conventional Optimization and Heuristic Approaches
Resource allocation in wireless networks has been addressed through formal mathematical optimization and heuristic algorithms [20, 21, 22, 25]. Algorithms such as Linear and Non-Linear Programming provide a rigorous framework for defining these problems, and for specific cases like convex problems, can even yield globally optimal solutions efficiently. When exact solutions are computationally infeasible due to the NP-hard nature of many practical scenarios involving discrete decisions, a wide range of heuristic and meta-heuristic algorithms, including greedy methods, simulated annealing, and genetic algorithms, have been employed. These methods aim to find good, albeit often suboptimal, solutions quickly, providing a practical alternative to complex optimization.
Despite their utility, these conventional approaches have significant limitations in the context of dynamic, sliced O-RAN environments. Their primary failing is the computational complexity that makes them too slow for the near-real-time control loops required by the O-RAN RIC. Furthermore, these methods typically rely on static models and lack the agility to adapt to the highly dynamic nature of wireless channels, traffic loads, and user mobility. Their performance is tied to the accuracy of underlying network models, and heuristics can be brittle, performing poorly when conditions deviate from their design assumptions. Crucially, they do not learn from operational data to improve future decisions, which makes them ill-suited for the data-rich, learning-driven paradigm enabled by O-RAN’s open interfaces.
3.2 Discriminative AI and Reinforcement Learning Paradigms
The limitations of conventional optimization have driven research into AI-driven solutions, with DRL emerging as a significant approach for dynamic resource allocation. DRL combines the function approximation capabilities of Deep Neural Network (DNN) with the trial-and-error framework of RL, enabling agents to learn optimal policies through interaction with their environment. Various DRL algorithms, including value-based methods like Deep Q-Network (DQN) and actor-critic methods such as Soft Actor-Critic, have been utilized in wireless resource management [26]. In complex systems like O-RAN, Multi-Agent Reinforcement Learning (MARL) allows multiple agents to learn simultaneously, optimizing local and global objectives. FL has also gained importance as a distributed machine learning paradigm that facilitates collaborative training without centralizing data, thus preserving privacy. The combination of DRL and FL, termed Federated Deep Reinforcement Learning (FDRL), allows distributed agents to enhance DRL model building. They offer advantages such as adaptability, reduced reliance on precise network models, the ability to manage high-dimensional state spaces, and enhanced data privacy [27]. However, DRL faces challenges such as sample inefficiency, long training times, and poor generalization in new scenarios. Additionally, designing practical reward functions is complex, and the ”black box” nature of deep networks raises interpretability concerns. Moreover, FL struggles with non-identically distributed data and communication overhead. DRL effectively learns reactive policies but does not model the generative processes or distributional characteristics of optimal solutions, highlighting the need to explore generative AI approaches to understand these underlying data distributions better.
3.3 The Rise of Generative AI in Wireless Network Optimization
A new frontier in AI-driven network optimization is emerging with the application of GAI. Unlike discriminative models that learn decision boundaries or RL agents that learn reactive policies, GAI models are designed to learn the underlying probability distribution of data. This allows them to generate new, synthetic data instances or solutions that exhibit similar characteristics to the training data [28]. The unique capabilities of GAI offer several advantages for complex optimization problems, such as O-RAN resource allocation, including the ability to proactively generate novel and optimized solutions, augment limited training datasets, and robustly model the inherent uncertainty of wireless environments by capturing complex data distributions [29].
This generative capability is realized through several prominent model architectures. VAEs are adept at learning compressed, meaningful representations of high-dimensional network data, making them suitable for both generating solutions and performing inference [1, 49]. Generative Adversarial Networks uses a competitive training process and has been primarily explored for data augmentation, creating realistic synthetic channel data or traffic patterns to enhance the robustness of other AI models [35, 45]. More recently, Diffusion Models have shown promise for generating high-fidelity samples and are being investigated for imitating expert allocation policies [30]. The choice of model often depends on the specific task, whether it is direct solution generation, data augmentation, or learning latent representations for other processes.
3.4 Identifying the Research Gap: Positioning the Proposed SS-VAE Framework
Despite notable advancements, current resource allocation methods in sliced O-RAN environments still face significant challenges. Conventional optimization and heuristic approaches often lack the speed required for real-time control and struggle to adapt to dynamic network conditions. DRL provides adaptability but suffers from poor sample efficiency and limited generalization, as it typically learns reactive policies without fully modeling the optimal solution distribution. Advanced techniques, such as FL and Transfer Learning (TL), offer some benefits but do not adequately address the need to learn from limited labeled data while exploring the vast solution space of NP-hard allocation problems.
There is a pressing need for approaches that can learn robust representations from scarce or mixed-quality data, model complex uncertainties, and explore the solution space in a generative manner to find near-optimal resource allocation strategies. Specifically, this is essential for coordinating multiple xApps within dynamic O-RAN environments. The challenge is heightened by the limited generalizability of existing models across diverse deployment conditions, including varying numbers of User Equipment (UE), fluctuating traffic patterns, and heterogeneous network configurations. This limits their scalability and effectiveness in real-world scenarios.
To address these issues, we propose the Generative Semi-Supervised VAE-Contrastive Learning (SS-VAE) framework, which integrates key AI concepts to fill this gap. At its core, the VAE learns the underlying probability distributions of network states and optimal resource allocations, enabling proactive generation and refinement of allocation decisions. This generative capability is vital for navigating the complex solution landscape of joint resource allocation, moving beyond the reactive tendencies of DRL and the data demands of purely supervised learning. Table I summarizes the key characteristics, contributions, and limitations of selected related works alongside the proposed SS-VAE approach.
TABLE I: Comparative Analysis of AI-Driven Resource Allocation Methods in O-RAN
Reference Problem Domain Methodology Key Contribution(s) Limitation(s) Primary Differentiator vs. This Work
SS-VAE Multi-resource (power, PRB, UE assoc.) allocation for eMBB/URLLC slicing in O-RAN. Semi-Supervised Generative AI (VAE with Contrastive Learning). Novel unified GAI framework. Data-efficient via semi-supervision. Generative exploration of solution space. Requires initial limited labels from a high-complexity solver. (Baseline for comparison)
Li et al. [31] Proactive resource allocation considering UE mobility in HetNets. DRL. Adapts proactively to user mobility. Standard DRL challenges (e.g., sample complexity, generalization). Learns a reactive policy, whereas our work is generative, learning the solution distribution.
Wu et al. [32] Joint radio and computation resource allocation for vehicular RAN slicing. DRL. Dynamically allocates both radio and compute resources. Relies on reactive DRL policy; generalization across diverse slice types. Reactive DRL policy vs. our generative exploration of the joint resource space.
Wang et al. [33] Resource management for O-RAN RU and DU entities. Self-Play DRL. Achieves adaptive decision-making through a self-play paradigm. Scalability and generalization of self-play in a complex O-RAN environment. Game-theoretic DRL policy vs. our GAI approach to learning underlying data characteristics.
Ndikumana et al. [34] Joint task offloading and fronthaul routing in O-RAN to reduce delay. FL + DRL. Ensures data privacy via FL while optimizing routing decisions. High complexity in coordinating FL and DRL; potential convergence issues with non-IID data. Augments a reactive DRL policy with FL, while our method is a unified generative model.
Mhatre et al. [35] Efficient resource management in 6G networks. TL + DRL. Reduces DRL training overhead by transferring knowledge from related tasks. TL is task-dependent; still relies on a reactive DRL core. Aids a reactive policy with TL; our method aims for broader generalization via distributional learning.
Chen & Heydari [36] Resource governance via adaptive adjustment of network topology. VAE + DRL. Optimizes network structure using an RL agent in a VAE-compressed latent space. High training complexity; uses VAE for state compression, not generation. VAE is used for state representation for an RL agent; our method is a unified GAI for direct solution generation.
Qiao et al. [37] Multi-timescale resource allocation for O-RAN slicing (radio & computing). Parallel Hierarchical DRL. Decomposes the problem hierarchically to handle complexity and coupled constraints. High complexity of the hierarchical DRL architecture. Advanced DRL structure that still learns reactive policies; our method is a single, unified generative model.
Lotfi et al. [26] Dynamic O-RAN slicing and resource management. LLM + MARL. Enriches RL state with semantic context from an LLM for improved planning. High computational cost of LLMs; relies on the quality of LLM output. Augments DRL state with LLM semantics; we learn robust features directly from raw data via contrastive learning.
Salama et al. [27] Resource allocation in O-RAN specifically to support FL processes. FL + RL + Model-based. Optimizes O-RAN resources (e.g., power) to improve the efficiency of an ongoing FL task. Solves a different, specific problem (RA for FL), not a general network slicing problem. Different problem domains and uses a hybrid of non-generative methods.
Qazzaz et al. [38] Dynamic PRB allocation in O-RAN based on traffic and QoS. Supervised ML (Random Forest Classifier). Provides fast decisions by selecting the best policy from a predefined, finite set of options. Cannot generate novel allocation policies; the quality of predefined policies bounds performance. A discriminative model that selects a policy; our model is generative and creates new allocation solutions.
4 System Model
This study addresses the resource allocation problem in the O-RAN architecture through network slicing, focusing on two primary slice types: URLLC and eMBB. The URLLC slice is designed for ultra-low latency and high-reliability communications, which are essential for critical applications. In contrast, the eMBB slice aims to deliver high data rates for mobile broadband services. Each slice type imposes distinct QoS requirements on resource allocation, as conceptually illustrated for the O-RAN functional split in Figure 1 [6]. Based on Figure 1, the O-RU is shared across services. Within the O-DU, the High PHY layer is common, while the MAC and RLC layers are logically separated per slice. In the O-CU-UP, the SDAP and PDCP layers are also sliced to cater to differentiated service needs.
[…]